OpenStudy (anonymous):

Please help Simplify the radical expression, use absolute value symbols when needed: ^5square root sign x^5 y^10

OpenStudy (anonymous):

y^10

OpenStudy (anonymous):

Or rather xy^10 I think

jhonyy9 (jhonyy9):

\[\sqrt[5] x^{5}\] = x

jhonyy9 (jhonyy9):

\[\sqrt[5]y^{10}\]

jhonyy9 (jhonyy9):

what will be equal

OpenStudy (anonymous):

\[\sqrt[5]{x^5} y^10\]

OpenStudy (anonymous):

please help

jhonyy9 (jhonyy9):

the result is xy2 what want signeing x and y squared